Secret Message - 2

Given a string, write a program to print a secret message that replaces characters with numbers 'a' with 1, 'b' with 2, ..., 'z' with 26 where characters are separated by '-'.




a b c d e f g h i j k l m n o p q r s t u v w x y z

1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26


Input


Input will be a string in single line containing spaces and letters both uppercase and lowercase.


Output


Output should be a single line containing secret message. All characters in output should be in lower case.


Explanation


For example, if input is "python", "p" should replaced with "16", similarly"y" with "25","t" with "20","h" with "8","o" with "15","n" with "14". So output be "16-25-20-8-15-14".


Sample Input 1

python


Sample Output 1

16-25-20-8-15-14




Sample Input 2

python learning


Sample Output 2

16-25-20-8-15-14 12-5-1-18-14-9-14-7
1
Expert's answer
2021-04-29T05:41:29-0400
def SecretMsg(Str):
    s = InputStr.lower()
    Msg=""
    for r in range(0,len(s)):
        if(ord(s[r])>=97 and ord(s[r])<=122):
           n = ord(s[r])-96
           if(r<len(s)-1):
                Msg = Msg+str(n)+"-"
           else:
                Msg = Msg+str(n)
    return(Msg)    


InputStr = "python"
print("\nSample Input: ",InputStr)
print("Output: ",SecretMsg(InputStr))


InputStr = "python learning"
print("\nSample Input: ",InputStr)
print("Output: ",SecretMsg(InputStr))